When Lee was 3 times as old as kevin. his sister Kate was 27. Kevin was half as old as Kate brother Lee was 38 there ages add to 143 how old are Kevin , Kate and Lee.

Answer by ankor@dixie-net.com(22740)

When Lee was 3 times as old as Kevin, his sister Kate was 27. Kevin was half as old as Kate brother Lee was 38 there ages add to 143 how old are Kevin , Kate, and Lee.
:
This is really poorly written. Whoever wrote this, should go back to pushing a broom.
:
Write an equation for each statement:
:
"When Lee was 3 times as old as kevin."
L = 3Ke
:
his sister Kate was 27.
Ka = 27
:
Rewriting the statement to make sense.
Kevin was half as old as Lee, when he was 38.
Ke = .5L = 38
Therefore:
L = 2 * 38 = 76
:
:Find out how many years ago (x) that L = 3Ke:
76 - x = 3(38-x)
76 - x = 114 - 3x
+3x - x = 114 - 76
2x = 38
x = 19 years ago
:
19 yrs ago L = 57; Ke = 19; (3*19=57); Ka = 27
Ka age now: 27 + 19 = 46 yrs old
:
Now the problems is that 38 + 76 + 46 = 160 not 143